40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 7 17:39:40 2012, keiko, Update, LSC, MICH noise budget on 5 Mar Mar5-MICHbudget.pngMar5-MICHbudget.png
    Reply  Wed Mar 7 20:53:13 2012, keiko, Update, LSC, MICH noise budget on 5 Mar Mar6-MICHbudget.png
    Reply  Thu Mar 8 00:57:48 2012, keiko, Update, LSC, MICH noise budget on Mar 5, Mar 6, and old Mar7MICHbudgettotal.png
Message ID: 6376     Entry time: Wed Mar 7 17:39:40 2012     Reply to this: 6380   6385
Author: keiko 
Type: Update 
Category: LSC 
Subject: MICH noise budget on 5 Mar 

 This is the calibrated MICH noise budget on Mar 5. There was a sharp peak at 1Hz and a blob on 3 Hz. The demod phase was adjusted for AS55Q.

Mar5-MICHbudget.png

 

Attachment 1: Mar5-MICHbudget.png  67 kB  | Hide | Hide all
Mar5-MICHbudget.png
ELOG V3.1.3-